OpenStudy (anonymous):

Evaluate [(51 + 3) - 32] ÷ 9 • 2.

OpenStudy (texaschic101):

[ (51 + 3) - 32 ] / 9 * 2 = [ 54 - 32 ] / 9 * 2 = [ 22] / 9 * 2 = 22/9 * 2 = 44/9 = 4 8/9

OpenStudy (anonymous):

Hey, welcome to Open Study! Your question is fine, but try posting in the 'math' subject next time. You can pick which subject to work on in the upper left hand side, called "Find More Subjects". Click on it, and search/find the subject your homework is. The reason you have to pick the right subject is because if you post in the English or Science section for History or Math, the answers can sometimes be irrelevant, and no one wants a bad grade from simply putting the question in the wrong subject! :) Thanks. :D